If it takes a train 7h to get from Capital City when it travels 40 mi/h,how long would it take the train to go the same distance when it travels 70 mi/h?Would it take 7/4 h,4h, 4/7 h or 10 h?
Answer by checkley75(3666)

distance=rate*time so we have d=7*40=280 miles.
280=70T
t=280/70
t=4 hours is the answer.